Problem 1. (20 points) Find the general solution to the following differential equations. (a) ty + 5ty - 5y = 0.
Answer #1

- Solution : 7/6 Given differential equation. ty + styl -sy=0 -- 0 Let t- ez z = logt. at = 4 lie. it yt) = yCz)). Peen plwing equation kW in equo, we gel - y (2) -y (2) + 5y (2) – 54 12) = 0 y(2) +4y(2) - 5y62) = 0 - © Auxillary eq m² +4 m -